# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

### v1

@@ III Data sources and Cooperative Perception in Autonomous Driving
In this section, we introduce the sensors and their modalities mainly used in autonomous driving (III-A). Furthermore, in III-B, we analyze the data acquisition and communication domain, such as onboard, drone, and vehicle-to-everything.
III-A Sensors and Modalities of Data
@@
Figure 4: Sensors on autonomous driving vehicle. The type of each sensor is (a) Camera: Basler ace acA1600-20uc, (b) LiDAR: Velodyne Puck LITE, (c) Radar: Ainstein Launches K-79, (d) Event-based camera: Evaluation Kit 4 HD, (e) IMU: IMU383_Aceinna-W and (f) Thermal camera: FLIR_2nd_Gen_ADK. All figures are extracted from the websites hosting the sensors.
Figure 5: Modalities of sensors. We demonstrate each type of sensor’s modality to intuitively understand each sensor’s characters.
Efficiently and precisely collecting data from the surrounding environment is the key to a reliable perception system in autonomous driving. To achieve this goal, various types of sensors are utilized on self-driving vehicles and infrastructures. The sensor examples are shown in Fig. 4. The most used sensors are cameras, LiDAR, and radars. Event-based and thermal cameras are also installed in vehicles or roadsides to improve the perception capability further.
RGB Images. RGB images are usually recorded by monocular, stereo, or fisheye cameras. Monocular cameras offer a 2D view without depth; stereo cameras, with their dual lenses, provide depth perception; fisheye cameras use wide-angle lenses to capture a broad view. All these cameras channel light through a lens onto an image sensor, e.g., CMOS, converting this light into electronic signals representing an image. As shown in Fig. 5 (a), the 2D images capture color information, rich textures, patterns, and visual details of the environment. Due to these characters, RGB images are mainly used to detect vehicles and pedestrians and recognize road signs. However, the RGB images are vulnerable to conditions like low illumination, rain, fog, or glare [43].
LiDAR Point Clouds. LiDAR uses laser beams to measure the distance between the sensor and an object, creating a 3D environment representation [44]. The LiDAR point clouds (Fig. 5 (b)) provide precise spatial information with high resolution and can detect objects over long distances. However, the density of these points can decrease with increasing distance, leading to sparser representations for distant objects. The weather conditions, e.g., fog, also limit the performance of LiDAR. In general, LiDAR is suitable in cases that require 3D concise information.
@@
Inertial Measurement Unit (IMU). An IMU is an electronic device that measures and reports an object’s specific force, angular rate, and sometimes magnetic field surrounding the object [48]. In autonomous driving, it is used to track the movement and orientation of the vehicle. Although IMU does not include visual information of the surrounding environment, the perception systems can achieve more accurate and robust tracking of a vehicle’s motion and orientation by fusing the data from an IMU with data from other sensors.
Figure 6: Sensor number distribution. We statistic the distribution of different sensors. Overall, RGB cameras and LiDAR are the most used sensors in autonomous driving datasets.
We analyze the sensor distribution from the collected datasets, shown in Fig. 6. More than half of the sensors are monocular cameras (53.85%) due to their low price and reliable performance. Additionally, 93 datasets include LiDAR data, valued for its high resolution and precise spatial information. However, its high-cost limits LiDAR’s widespread use. Beyond LiDAR point clouds, 29 datasets leverage stereo cameras to capture depth information. Furthermore, 5.41%, 3.42%, and 1.71% datasets include radar, thermal camera, and fisheye camera, respectively. Given the temporal efficiency of capturing dynamic scenes, three datasets generate data based on event-based cameras.
@@ III-B Sensing Domains and Cooperative Perception Systems
The sensory data and communication between the ego vehicle and other entities in the surrounding environment play a pivotal role in ensuring self-driving systems’ safety, efficiency, and overall functionality. Therefore, the positioning of sensors is crucial as it determines the quality, angle, and scope of data that can be collected. Generally, sensors in the autonomous driving environment can be categorized into the following domains: onboard, Vehicle-to-Everything (V2X), drone, and others.
Figure 7: Overview of cooperative perception systems in autonomous driving. A complete autonomous driving perception system consists of the ego vehicle and the cooperation between vehicles, infrastructure, and networks.
Onboard. Onboard sensors are installed directly on the autonomous driving vehicle and usually consist of cameras, LiDAR, radars, and IMUs. These sensors provide a direct perspective from the vehicle’s standpoint, offering immediate feedback on the surroundings. Nevertheless, due to the limitation of the vehicle detection scope, onboard sensors may have limitations in providing advanced warning about obstacles in blind spots or detecting hazards around sharp bends.
V2X. Vehicle-to-Everything (V2X) encompasses communications between a vehicle and any other components in the transport system [49], including V2V, V2I, and V2N (as shown in Fig. 7). Beyond the immediate sensory input, the cooperative systems ensure multiple entities work harmoniously.
@@
    V2V enables nearby vehicles to share data, including their positions and velocity and sensory data, like camera images or LiDAR scans, which assists in creating a more comprehensive understanding of the driving scenarios.
@@
    V2I facilitates communications between the autonomous vehicle and infrastructure components such as traffic lights, signs, or roadside sensors. The sensors embedded in the road infrastructure, including cameras, LiDAR, radars, or event-based cameras, work synergistically to extend the perception range and improve the situational awareness of autonomous vehicles. In this survey, we consider both perceptions through infrastructure or V2I belonging to V2I.
@@
Drone. Drones, or UAVs, offer an aerial perspective, providing data essential for trajectory prediction and route planning [18] . For example, the real-time data from drones can be integrated into traffic management systems to optimize the traffic flow and alert autonomous vehicles of accidents ahead.
@@ IV Tasks in Autonomous Driving
This section offers insight into the pivotal tasks in autonomous driving, such as perception and localization (IV-A), prediction (IV-B), and planning and control (IV-C). The overview of the autonomous driving pipeline is demonstrated in Fig. 8. We detail their objectives, the nature of data they rely upon, and inherent challenges. Fig. 9 illustrates examples of several main tasks in autonomous driving.
Figure 8: The overview of autonomous driving pipeline. Autonomous driving systems can be categorized into two types: modular-based and end-to-end. Both rely on the data collected by various sensors installed on the vehicles or infrastructures and interact with the surrounding driving experiments.
@@ IV-A Perception and Localization
Figure 9: Examples of various autonomous driving tasks. (a) is from KITTI [13], (b) is from Cityscapes [14], (c) is from V2X-Seq [50], (d) is from BDD100K [24], (e) is from Refer-KITTI [51], (f) is from KITTI-360 [52], (g) is from Dr(eye)ve [53], (h) is from Argoverse 2 [28]. All figures are collected from the open-source data of datasets or the websites hosting the datasets.
@@
2D/3D Object Detection. 2D or 3D object detection aims to identify and classify other entities within the driving environment. While 2D object detection identifies objects in image space, 3D object detection further incorporates precise depth information, often provided by LiDAR. Although the detection technologies have significantly advanced, several challenges remain, such as object occlusions, varying light conditions, and diverse object appearances.
@@
$mIoU=\frac{1}{k=1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}+\sum_{j=0}^{k}p_{ji}-p_{ii}}$ (8)
@@
HD Map. HD mapping aims to construct detailed, highly accurate representations that include information about road structures, traffic signs, and landmarks. A dataset should provide LiDAR data for precise spatial information and camera data for visual details to ensure established map accuracy. According to [28], HD map automation [58] and HD map change detection [59] have received more and more attention. Normally, the HD map quality is estimated using the accuracy metric.
SLAM. Simultaneous Localization And Mapping (SLAM) entails building a concurrent map of the surrounding environment and localizing the vehicle within this map. Hence, data from cameras, IMUs for position tracking, and real-time LiDAR point clouds are vital. [60] introduces two evaluation metrics, relative pose error (RPE) and absolute trajectory error (ATE), for evaluating the quality of the estimated trajectory from the input RGB-D images.
@@ IV-B Prediction
Trajectory Prediction. Using time-series data from sensors like cameras and LiDAR, trajectory prediction pertains to anticipating the future paths or movement patterns of other entities [61], such as pedestrians, cyclists, or other vehicles.
Behavior Prediction. Behavior prediction anticipates the potential actions of other road users [62], e.g., whether a vehicle will change lanes. Training behavior prediction models rely on extensive annotated data due to entities’ vast range of potential actions within various scenarios.
Intention Prediction. Intention prediction focuses on inferring the higher-level goals of the intention behind the actions of objects, involving a deeper semantic comprehension of the physical or mental activities of humans [64]. Because of the task’s complexity, it requires data not only from perception sensors like cameras but also additional information, such as traffic signals and hand gestures, to infer the intentions of other agents.
@@ IV-C1 Planning
Planning represents the decision-making process in reaction to the perceived environment and predictions. A classic three-level hierarchical planning framework consists of path, behavioral, and motion planning [53].
@@ IV-D End-to-end Autonomous Driving
End-to-end approaches in autonomous driving are those where a single deep learning model handles everything from perception to control, bypassing the traditional modular pipeline. Such models can often be more adaptive as they rely on adjusting the whole model through learning. Their inherent promise is in their simplicity and efficiency by reducing the need for hand-crafted components [66]. However, implementing end-to-end models faces critical limitations such as extensive training data requirements, low interpretability, and inflexible modular tuning.
@@ V High-Influence Datasets
Figure 10: Chronological overview. We show the top 50 datasets with the highest impact scores. This diagram records datasets from 2007 to 2023 (until the writing of this paper).
Dataset Year Size Temp Tasks Categories Weather Time of day Scenario Geometrical Impact
@@
nuScenes [22] 2019 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 3 2 4 2 81.20
Cityscapes [14] 2016 25K $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 30 1 1 3 1 79.00
BDD100K [24] 2020 12M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 40 5 2 4 1 77.82
Waymo [23] 2019 230K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 2 3 5 1 77.81
KITTI [13] 2012 41K $\checkmark$ $\checkmark$ $\checkmark$ 8 1 1 2 1 77.30
SYNTHIA [2] 2016 13.4K $\times$ $\checkmark$ 13 2 3 5 - 70.44
Apolloscape [17] 2018 143,906 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 28 4 3 3 1 68.97
Virtual KITTI [3] 2016 21,260 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 3 2 5 - 67.61
VIPER [5] 2017 254,064 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11 4 3 4 - 67.07
SemanticKITTI [6] 2019 43,552 $\checkmark$ $\checkmark$ 28 1 1 4 1 66.66
GTA5 [4] 2016 24,966 $\times$ $\checkmark$ 19 2 2 2 - 66.10
Argoverse 2 [28] 2023 6M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30 2 1 1 6 61.11
Lane Det [68] 2017 133,235 $\times$ $\checkmark$ 1 1 3 3 1 59.62
CityPersons [69] 2017 25K $\times$ $\checkmark$ $\checkmark$ 30 2 1 - 27 58.84
CamVid [70] 2009 701 $\times$ $\checkmark$ 32 2 1 2 1 58.69
IDD [71] 2019 10,004 $\times$ $\checkmark$ 34 3 3 5 1 58.12
Foggy Cityscapes [72] 2018 20,550 $\times$ $\checkmark$ $\checkmark$ 19 1 1 2 1 58.09
A2D2 [25] 2020 41,277 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 38 2 1 3 3 58.06
Virtual KITTI 2 [9] 2020 20,992 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 4 2 3 1 56.74
RADIATE [26] 2021 44,140 $\checkmark$ $\checkmark$ $\checkmark$ 8 5 2 4 1 56.70
GTSDB [73] 2013 900 $\times$ $\checkmark$ 4 2 2 3 1 56.68
Caltech Ped [74] 2009 250K $\times$ $\checkmark$ 1 1 1 1 2 56.64
SHIFT [10] 2022 2,5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 5 5 3 - 55.79
CAOS [7] 2022 13K $\times$ $\checkmark$ 13 3 2 3 - 55.74
STF [75] 2020 13,500 $\times$ $\checkmark$ $\checkmark$ 1 4 2 3 4 54.91
KITTI-360 [52] 2021 150K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37 1 1 1 1 54.77
ACDC [76] 2021 4,006 $\times$ $\checkmark$ 19 4 2 3 1 54.62
VPGNet [77] 2017 20K $\times$ $\checkmark$ 17 3 2 1 1 54.37
ONCE [29] 2021 1M $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 5 3 2 5 1 54.09
@@
DAIR-V2X [27] 2021 71,254 $\checkmark$ $\checkmark$ $\checkmark$ 10 2 2 2 1 49.21
V2XSet [11] 2022 11,447 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 1 - 45.39
V2VNet [78] 2020 51,2K $\checkmark$ $\checkmark$ 1 1 1 1 - 44.94
T&J [79] 2019 100 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 2 1 44.58
Rope3D [80] 2022 50K $\times$ $\checkmark$ $\checkmark$ 12 3 3 2 1 44.39
V2X-Sim [81] 2022 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 1 1 1 - 41.00
@@
Co-Percep [83] 2020 10K $\times$ $\checkmark$ $\checkmark$ 1 1 1 1 1 40.06
A9-Dataset [84] 2022 1,098 $\checkmark$ $\checkmark$ $\checkmark$ 9 4 2 1 1 39.12
LUMPI [85] 2022 200K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 3 6 3 1 1 35.87
DOLPHINS [86] 2022 42,376 $\checkmark$ $\checkmark$ $\checkmark$ 2 3 2 4 - 35.21
@@
TABLE II: High-impact perception datasets. For a more comprehensive demonstration, we exhibit 50 perception datasets from different sensing domains instead of those with the highest scores.

### v2

@@ III Sensors and Perception Technology in Autonomous Driving
In this section, we introduce the sensors and their modalities mainly used in autonomous driving (III-A). Subsequently, in III-B, we analyze the data acquisition methods and cooperative perception technologies.
III-A Sensor Modalities
@@
Figure 4: Sensors on autonomous driving vehicles. The type of each sensor is (a) Camera: Basler ace acA1600-20uc, (b) LiDAR: Velodyne Puck LITE, (c) Radar: Ainstein Launches K-79, (d) Event-based camera: Evaluation Kit 4 HD, (e) IMU: IMU383_Aceinna-W and (f) Thermal camera: FLIR_2nd_Gen_ADK. All figures are extracted from the websites hosting the sensors.
Figure 5: We present the sensor modalities to provide an intuitive understanding of each sensor’s characteristics. (a) is from nuScenes [22], (b) is from KITTI [13], (c) is from [46], (d) is from [47], (e) is from [48]. All figures are collected from the open-source data of datasets.
Efficiently and precisely collecting data from the surrounding environment is the key to a reliable perception system in autonomous driving. To achieve this goal, various types of sensors are utilized on self-driving vehicles and infrastructures. The sensor examples are shown in Fig. 4. The most used sensors are cameras, LiDARs, and radars. Event-based and thermal cameras are also installed on vehicles or roadside infrastructure to improve the perception capability further.
RGB Images. RGB images are usually recorded by monocular, stereo, or fisheye cameras. Monocular cameras offer a 2D view without depth; stereo cameras, with their dual lenses, provide depth perception; fisheye cameras use wide-angle lenses to capture a broad view. As shown in Fig. 5 (a), the 2D images capture color information, rich textures, patterns, and visual details of the environment. Due to these characters, RGB images are mainly used to detect vehicles and pedestrians and recognize road signs. However, the RGB images are vulnerable to conditions like low illumination, rain, fog, or glare [49].
LiDAR Point Clouds. LiDARs use laser beams to measure the distance between the sensor and an object, creating a 3D environment representation [50]. LiDAR point clouds (Fig. 5 (b)) provide precise spatial information with high resolution and can detect objects over long distances. However, the density of these points can decrease with increasing distance, leading to sparser representations for distant objects. Weather conditions, e.g., fog, also limit the performance of LiDAR. In general, LiDARs are suitable in cases that require 3D concise information.
@@
Inertial Measurement Unit (IMU). An IMU is an electronic device that measures and reports an object’s specific force, angular rate, and sometimes magnetic field surrounding the object [54]. In autonomous driving, it is used to track the movement and orientation of the vehicle.
We analyze the sensor distribution from the collected datasets, shown in Fig. 6. More than half of the sensors are monocular cameras (52.79%) due to their low price and reliable performance. Additionally, 93 datasets (25.98%) include LiDAR data, valued for its high resolution and precise spatial information. However, its high-cost limits LiDAR’s widespread use. Beyond LiDAR point clouds, 29 datasets leverage stereo cameras to capture depth information. Furthermore, 5.31%, 3.35%, and 1.68% datasets include radar, thermal camera, and fisheye camera. Given the temporal efficiency of capturing dynamic scenes, ten datasets generate data based on event-based cameras (2.79%).
Figure 6: Sensor type distribution. We show the distribution of different sensors. Overall, RGB cameras and LiDARs are the most used sensors in autonomous driving datasets.
@@ III-B Sensing Domains and Cooperative Perception Systems
Sensory data and cooperation between the ego vehicle and surrounding entities are pivotal for ensuring autonomous driving systems’ safety, efficiency, and overall functionality. Therefore, the positioning of sensors is crucial as it determines the quality, angle, and scope of data that can be collected. Generally, sensing domains in autonomous driving can be categorized as onboard, Vehicle-to-Everything (V2X), drone-based, and others.
Figure 7: Overview of cooperative perception systems in autonomous driving. A complete autonomous driving perception system consists of the ego vehicle along with collaborative interactions between vehicles, infrastructure, and networks.
Onboard. Onboard sensors are installed directly on the autonomous vehicle and usually consist of cameras, LiDARs, radars, and IMUs. These sensors provide a direct perspective from the vehicle’s standpoint, offering immediate feedback on the surroundings. Nevertheless, due to the limited detection scope, onboard sensors may have limitations in providing advanced warnings about obstacles in blind spots or detecting hazards around sharp bends.
V2X. Vehicle-to-Everything encompasses communications between a vehicle and any other components in the transport system [55], including Vehicle-to-Vehicle, Vehicle-to-Infrastructure, and Vehicle-to-Network (as shown in Fig. 7). Beyond the immediate sensory input, the cooperative systems ensure multiple entities work harmoniously.
@@
    V2V enables nearby vehicles to share essential data, such as position, velocity, and captured sensory data (e.g., camera images or LiDAR scans). This shared information contributes to a more comprehensive understanding of driving scenarios.
@@
    V2I facilitates communications between the autonomous vehicle and infrastructure components such as traffic lights, signs, or roadside sensors. The sensors implemented in road infrastructure collaborate to enhance the perception range and situational awareness of autonomous vehicles. In this survey, we categorize interactions involving single or multiple vehicles with single or multiple infrastructure components, as well as collaborations among multiple infrastructure elements, under V2I.
@@
Drone. Drones, or Unmanned Aerial Vehicles (UAVs), offer an aerial perspective, providing data essential for trajectory prediction and route planning [18]. For example, the real-time data from drones can be integrated into traffic management systems to optimize the traffic flow and alert autonomous vehicles of accidents ahead.
@@ IV Tasks in Autonomous Driving
This section offers insight into the pivotal tasks in autonomous driving, such as perception and localization (IV-A), prediction (IV-B), and planning and control (IV-C). The overview of the autonomous driving pipeline is demonstrated in Fig. 8. We detail their objectives, the nature of data they rely upon, and inherent challenges. Fig. 9 highlights several main tasks in autonomous driving.
Figure 8: The overview of autonomous driving pipeline. Autonomous driving systems can be categorized into two types: modular-based and end-to-end. Both types rely on data collected by various sensors installed on vehicles or infrastructures. These systems interact with and respond to the surrounding environment during driving scenarios.
@@ IV-A Perception and Localization
Figure 9: Examples of various autonomous driving tasks. (a) is from KITTI [13], (b) is from Cityscapes [14], (c) is from V2X-Seq [56], (d) is from BDD100K [24], (e) is from Refer-KITTI [57], (f) is from KITTI-360 [58], (g) is from Dr(eye)ve [59], (h) is from TUMTraf [60]. All figures are collected from the open-source data of datasets or the websites hosting the datasets.
@@
2D/3D Object Detection. 2D or 3D object detection aims to identify the locations and classification of other entities within the driving environment. Although the detection technologies have significantly advanced, several challenges remain, such as object occlusions, varying light conditions, and diverse object appearances.
@@
$mIoU=\frac{1}{k+1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}+\sum_{j=0}^{k}p_{ji}-p_{ii}}$ (8)
@@
HD Map. HD mapping aims to construct detailed, highly accurate representations that include information about road structures, traffic signs, and landmarks. A dataset should provide LiDAR data for precise spatial information and camera data for visual details to ensure established map accuracy. According to [28], HD map automation [65] and HD map change detection [66] have received more and more attention. Usually, the HD map quality is estimated using the accuracy metric.
SLAM. Simultaneous Localization And Mapping (SLAM) entails building a concurrent map of the surrounding environment and localizing the vehicle within this map. Hence, data from cameras, IMUs for position tracking, and real-time LiDAR point clouds are vital. Sturm et al. [67] introduces two evaluation metrics, relative pose error (RPE) and absolute trajectory error (ATE), for evaluating the quality of the estimated trajectory from the input RGB-D images.
@@ IV-B Prediction
Trajectory Prediction. Based on time-series data from sensors like cameras and LiDARs, trajectory prediction pertains to anticipating the future paths or movement patterns of other entities [68], such as pedestrians, cyclists, or other vehicles.
Behavior Prediction. Behavior prediction anticipates the potential actions of other road users [69], e.g., whether a vehicle will change the lane. Training behavior prediction models rely on extensive annotated data due to entities’ vast range of potential actions within various scenarios.
Intention Prediction. Intention prediction focuses on inferring the higher-level goals behind the actions of objects, involving a deeper semantic comprehension of the physical or mental activities of humans [71]. Because of the task’s complexity, it requires data from perception sensors like cameras, traffic signals, and hand gestures.
@@ IV-C1 Planning
Planning represents the decision-making process in reaction to the perceived environment and predictions. A classic three-level hierarchical planning framework comprises path, behavioral, and motion planning [59].
@@ IV-D End-to-end Autonomous Driving
End-to-end approaches in autonomous driving are those where a single deep learning model handles everything from perception to control, bypassing the traditional modular pipeline. Such models can often be more adaptive as they rely on adjusting the whole model through learning. Their inherent promise is in their simplicity and efficiency by reducing the need for hand-crafted components [73]. However, implementing end-to-end models faces limitations such as extensive training data requirements, low interpretability, and inflexible modular tuning.
@@ V High-Influence Datasets
Figure 10: Chronological overview. We show the top 50 datasets ranked by their impact scores. This diagram covers datasets released from 2009 through March 17, 2024.
TABLE II: High-impact perception datasets. For a more comprehensive demonstration, we exhibit 50 perception datasets from different sensing domains instead of those with the highest scores. The citation number of each dataset is not shown in the table due to the fast change for this point.
Dataset Year Size Temp Tasks Categories Weather Time of day Scenario Geographical Impact
@@
nuScenes [22] 2019 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 3 2 4 2 83.03
Cityscapes [14] 2016 25K $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 30 1 1 3 1 79.87
BDD100K [24] 2020 12M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 40 5 2 4 1 79.79
Waymo [23] 2019 230K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 2 3 5 1 79.66
KITTI [13] 2012 41K $\checkmark$ $\checkmark$ $\checkmark$ 8 1 1 2 1 78.04
SYNTHIA [2] 2016 13.4K $\times$ $\checkmark$ 13 2 3 5 - 71.16
Apolloscape [17] 2018 143,906 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 28 4 3 3 1 69.91
SemanticKITTI [6] 2019 43,552 $\checkmark$ $\checkmark$ 28 1 1 4 1 68.40
Virtual KITTI [3] 2016 21,260 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 3 2 5 - 68.27
VIPER [5] 2017 254,064 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11 4 3 4 - 67.96
GTA5 [4] 2016 24,966 $\times$ $\checkmark$ 19 2 2 2 - 66.87
Argoverse 2 [28] 2023 6M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30 2 1 1 6 65.28
Lane Det [75] 2017 133,235 $\times$ $\checkmark$ 1 1 3 3 1 60.71
SHIFT [10] 2022 2,5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 5 5 3 - 60.67
CityPersons [76] 2017 25K $\times$ $\checkmark$ $\checkmark$ 30 2 1 - 27 59.75
A2D2 [25] 2020 41,277 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 38 2 1 3 3 59.42
Foggy Cityscapes [77] 2018 20,550 $\times$ $\checkmark$ $\checkmark$ 19 1 1 2 1 59.38
CamVid [78] 2009 701 $\times$ $\checkmark$ 32 2 1 2 1 59.27
IDD [79] 2019 10,004 $\times$ $\checkmark$ 34 3 3 5 1 59.18
CAOS [7] 2022 13K $\times$ $\checkmark$ 13 3 2 3 - 58.98
RADIATE [26] 2021 44,140 $\checkmark$ $\checkmark$ $\checkmark$ 8 5 2 4 1 58.78
Virtual KITTI 2 [9] 2020 20,992 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 4 2 3 1 58.67
KITTI-360 [58] 2021 150K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37 1 1 1 1 58.48
Dr(eye)ve [59] 2018 555000 $\times$ $\checkmark$ 10 3 3 3 - 58.13
ACDC [80] 2021 4,006 $\times$ $\checkmark$ 19 4 2 3 1 57.64
GTSDB [81] 2013 900 $\times$ $\checkmark$ 4 2 2 3 1 57.48
ONCE [30] 2021 1M $\times$ $\checkmark$ $\checkmark$ 5 3 2 5 1 57.03
Caltech Ped [82] 2009 250K $\times$ $\checkmark$ 1 1 1 1 2 56.98
STF [83] 2020 13,500 $\times$ $\checkmark$ $\checkmark$ 1 4 2 3 4 56.85
@@
TUMTraf [84, 85, 86, 60] 2022 50,253 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 10 6 5 10 1 55.51
DAIR-V2X [27] 2021 71,254 $\checkmark$ $\checkmark$ $\checkmark$ 10 2 2 2 1 54.58
V2XSet [11] 2022 11,447 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 1 - 49.84
@@
Rope3D [88] 2022 50K $\times$ $\checkmark$ $\checkmark$ 12 3 3 2 1 48.96
V2X-Sim [89] 2022 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 1 1 1 - 48.50
V2VNet [90] 2020 51,2K $\checkmark$ $\checkmark$ 1 1 1 1 - 48.20
T&J [91] 2019 100 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 2 1 46.63
Co-Percep [92] 2020 10K $\times$ $\checkmark$ $\checkmark$ 1 1 1 1 1 42.99
DeepAccident [12] 2023 285K $\checkmark$ 6 4 3 2 - 41.79
LUMPI [93] 2022 200K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 3 6 3 1 1 40.42